A patient is started on furosemide. Lasix 2,000 mg in 200 ml is prescribed at 1 mg per min. You should set the IV pump at how many ml per hour?
Answer #1

ANSWER.

* Calculation of ml/hr.

        ml/hr = Volume(ml) / Time(hr).

    Calculation of volume(ml).

         Volume(ml) = Desired dose(D)/Dose in hand(H) *Quantity(Q)

   Here

      Desired dose(D) = 1 mg

      Dose in hand(H) = 2000 mg.

         Quantity(Q) = 200 ml;

As per above formula

     Volume(ml) of drug required = D/H* Q

                                                   = 1/2000 * 200

                                                  = 0.0005 * 200

                                                  = 0.1 ml.

                        Time = 1 min.

       Convert ml/min to ml/hr by multiply with 60

                        ml/hr = 0.1 ml/min * 60

                                 = 0.1 * 60

                                 = 6 ml/hr.